OpenStudy (anonymous):

Look at the formula for finding distance. d = rt Which of the following is the formula to find time (t)?

OpenStudy (anonymous):

t=d/r t=d-r t=r/d t=dr

OpenStudy (anonymous):

Okay, lets do this, first take r to the other side to isolate t,

OpenStudy (anonymous):

so the inverse of multiplication is division

OpenStudy (anonymous):

???

OpenStudy (anonymous):

t=d/r because we divide both sides by r to isolate t

OpenStudy (anonymous):

^You're supposed to walk them through it

OpenStudy (anonymous):

Nah its good

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i just did bro u are confusing him with all the long terms so calm down